Dublin University Rifle Club

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Dublin University Rifle Club Mark Dennehy 167 Applewood Hts Greystones Co. Wicklow Ireland

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

A Chara,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event and open only to NTSA members), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

For shooters from N.Ireland, DURC rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use (and for shooters from other clubs as well if necessary). Please contact me for further details on this.

Misé le meas, Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

University College Dublin Rifle Club UCD Rifle Club PO Box 70 Students Forum UCD Belfield Dublin 4

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

A Chara,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event and open only to NTSA members), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

For shooters from N.Ireland, DURC rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use (and for shooters from other clubs as well if necessary). Please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Fassaroe Shooting Club Dave Cooney Ardnaglass Corrig Road Dun Laoghaire

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

A Chara,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event and open only to NTSA members), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

For shooters from N.Ireland, DURC rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use (and for shooters from other clubs as well if necessary). Please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Eastern Command Rifle Club Dave Dunne 17 Templeview Place Clare Hall Dublin 17

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

A Chara,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event and open only to NTSA members), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

For shooters from N.Ireland, DURC rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use (and for shooters from other clubs as well if necessary). Please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Dublin Rifle Club John K. Walsh Carjon Lodge Cross Chapel Blessington Co. Wicklow

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

A Chara,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event and open only to NTSA members), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

For shooters from N.Ireland, DURC rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use (and for shooters from other clubs as well if necessary). Please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Comber Rifle Club Agnes Graham 114 Towerview Crescent Bangor Co. Down BT19 6A2

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

A Chara,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event and open only to NTSA members), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

For shooters from N.Ireland, DURC rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use (and for shooters from other clubs as well if necessary). Please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

East Antrim Rifle Club Thomas McKnight East Antrim Rifle & Pistol Club 26 Meadow Drive Ballyclare County Antrim BT39 9XF Northern Ireland

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

A Chara,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event and open only to NTSA members), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

For shooters from N.Ireland, DURC rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use (and for shooters from other clubs as well if necessary). Please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Rathdrum Rifle and Pistol Club Karen Farrell 41 Deerpark Walk Blessington Co. Wicklow

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

A Chara,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event and open only to NTSA members), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

For shooters from N.Ireland, DURC rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use (and for shooters from other clubs as well if necessary). Please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Wilkinstown Rifle Club Geoff Cooney Wilkinstown Navan Co. Meath

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

A Chara,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event and open only to NTSA members), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

For shooters from N.Ireland, DURC rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use (and for shooters from other clubs as well if necessary). Please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Condor Rifle Club Terry Wearen Duleek House Duleek Co. Meath

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

A Chara,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event and open only to NTSA members), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

For shooters from N.Ireland, DURC rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use (and for shooters from other clubs as well if necessary). Please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Fermoy Rifle Club Kieran Barry Coolmakee Fermoy Co. Cork

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

A Chara,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event and open only to NTSA members), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

For shooters from N.Ireland, DURC rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use (and for shooters from other clubs as well if necessary). Please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Fingal Rifle Club John Walls Conel Hill Balbriggan Co. Dublin

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

A Chara,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event and open only to NTSA members), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

For shooters from N.Ireland, DURC rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use (and for shooters from other clubs as well if necessary). Please contact me for further details on this.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Dungannon Rifle Club Robert McVeigh 86 Gilford Road Lurgan Co. Armagh BT66 7EB

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC 50m Open  17th July 2001

A Chara,

The DURC 50m Open is to be held this Sunday (July 22) in the Fassaroe Sporting Club range in Enniskerry.

First prize in the Championship is a voucher for a Shilen barrel blank in the winner’s choice of calibre and twist.

There will be the usual English Match in the morning (this is a Registered event and open only to NTSA members), followed by a barbecue and then a 50m Rapid-Fire event. This will be ten shots in ninty seconds, five shots per target. This is an annual competition at 25yards for DURC, and usually yields interesting results, the winner often being a surprise.

First detail kicks off at 0930, last entries are at 1230 and the final detail ends at 1400. Allowing a short break for the BBQ lunch, the 50m Rapid-Fire competition begins at 1500, and should be finished by 1600.

Entry Fees are £15 per entrant, which covers the English Match, NTSA levy, Rapid-Fire competition and the BBQ.

For shooters from N.Ireland, DURC rifles (all Anschutz 18-series and 19-series rifles with decent barrels) will be available for use (and for shooters from other clubs as well if necessary). Please contact me for further details on this.